# Leptogenesis from magnetic helicity of gauged $\text{U}(1)_{B-L}$

### v1

@@ III Model with Majorana neutrinos
In this section, we examine the dynamics of the $B-L$ charge in the fermion sector. We focus on the seesaw model Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979), in which the interaction between the right-handed neutrinos and the standard model particles results in the standard model neutrinos behaving as Majorana particles at low energies. We investigate if our scenario can account for the generation of the baryon asymmetry of the universe.
@@
Figure 2: Schematic picture of a successful scenario. First, the $B-L$ helicity (the blue line) is generated by, for example, the axion inflation, together with the $B-L$ charges in the standard model sector (the green line) and the chiral sector (the orange line). Then, the $B-L$ charges are (partially) equilibrated by the interactions between these two sectors to wash out the $B-L$ charge in the fermion sector. Finally, the $B-L$ helicity decays and the $B-L$ charges are generated in both sectors. The final $B-L$ charge in the standard model sector corresponds to the baryon number of the universe today.
@@
Let us first recapitulate our leptogenesis scenario and summarize the requirements for the model. We show a schematic picture of our scenario in Fig. 2. First, we assume the $B-L$ helicity is generated in the early universe. There are several mechanisms to generate the $B-L$ helicity, such as the axion inflation Turner and Widrow (1988); Garretson et al. (1992); Anber and Sorbo (2006); Barnaby and Peloso (2011); Barnaby et al. (2011) and the chiral plasma instability Joyce and Shaposhnikov (1997); Akamatsu and Yamamoto (2013); Rogachevskii et al. (2017); Schober et al. (2018); Kamada (2018). As we have discussed in Sec. II, we assume the axion inflation as the generation mechanism in this paper. When the $B-L$ helicity is generated, the $B-L$ charges in the standard model sector are also generated according to the anomaly equation, Eq. (II.2).In the case of chiral plasma instability Joyce and Shaposhnikov (1997); Akamatsu and Yamamoto (2013); Rogachevskii et al. (2017); Schober et al. (2018); Kamada (2018), we need to start from a chiral charge for the $B-L$ gauge interaction in the fermion sector. The distribution of $B-L$ charge in the standard model and chiral sector after $B-L$ helicity generation depends on the initial condition. As long as the $B-L$ symmetry is gauged, regardless of the detail of the model, some chiral sector is added to cancel the ’t Hooft anomaly of the $B-L$ symmetry in the standard model sector. The $B-L$ charge is therefore generated in the chiral sector as well. Note that the total $B-L$ charge of the universe is zero; the $B-L$ charges in the standard model sector and the chiral sector are the same amount with the opposite sign.
@@
As we have discussed in Sec. II, the magnetic helicity of the $B-L$ gauge field decays after the $B-L$ gauge symmetry is Higgsed. The decay generates the $B-L$ charge both in the standard model sector and the chiral sector. If the $B-L$ charges in both sectors are conserved separately from the beginning, the final $B-L$ charges in both sectors are zero. This can be seen from the integrated form of the anomaly equation, Eq. (II.3); the initial $B-L$ charges in both sectors and the magnetic helicity are zero and the final helicity is zero, so that the final $B-L$ charges in both sectors are zero.
@@
To generate non-zero $B-L$, we need two requirements. First, we need to introduce the interactions between the standard model particles and the chiral sector to equilibrate, at least partially, the $B-L$ charges in both sectors, as explained above. However, if this interaction does not decouple even after the decay of the magnetic helicity, the $B-L$ charges generated by the helicity decay in both sectors are also equilibrated, and the final $B-L$ charges in both sectors are zero. Thus, secondly, it is also required to decouple the interactions before the decay of the magnetic helicity. Then, the decay of the magnetic helicity generates the difference between the $B-L$ charges in the standard model sector and the chiral sector, which remains even today as the baryon number of the universe.
@@
Before discussing the details of the interactions, let us make a few comments on the above summary. Firstly, we have not specified the chiral sector yet. For the seesaw model discussed in this section, the chiral sector is the right-handed neutrinos before the $B-L$ symmetry breaking. However, as the right-handed neutrinos become extremely massive after the symmetry breaking, one may wonder what happens after the decoupling of the right-handed neutrinos. Once the right-handed neutrinos become massive, the $B-L$ symmetry is spontaneously broken and there is a Goldstone boson, $\phi_{M}$, although the Goldstone boson is eventually eaten by the $B-L$ gauge field. The Goldstone boson has the Wess-Zumino-Witten term and cancels the ’t Hooft anomaly Green and Schwarz (1984); Faddeev and Shatashvili (1986); Krasnikov (1985); Babelon et al. (1986); Harada and Tsutsui (1987); Preskill (1991). In other words, the Goldstone boson is chiral and plays the role of the chiral sector. The total $B-L$ charge is always zero even today if we include the unphysical contribution from $\phi_{M}$, as $B-L$ symmetry is gauged.
@@
Secondly, from the integrated anomaly equation, Eq. (II.3), one may wonder if the $B-L$ charge in the standard model sector is always zero given that the initial $B-L$ charge and initial and final helicity is vanishing. Contrary to this naive expectation, the final $B-L$ charge in the standard model sector is not zero. This is because the interactions equilibrating the $B-L$ charges in the standard model sector and the chiral sector explicitly respect only the total $B-L$ symmetry and break the $B-L$ symmetry in the standard model sector. Schematically, the total derivative of the standard model $B-L$ current is modified as
[…]
@@
[…]
$\mathcal{L}_{\Phi}$ |$\equiv\absolutevalue{\quantity(\partial_{\mu}-iq_{\Phi}g_{B-L}X_{\mu})\Phi}^{2}-V\quantity(|\Phi|^{2})$ (III.62)
@@
where $N_{i}$ is $i$-th ($i=1,2,3$) generation of the right-handed neutrino ******Readers should not confuse the name, “right-handed”, with the chirality of the right-handed neutrino. Throughout the paper, we use left-handed Weyl spinors for the chiral fermions., $\Phi$ is the Higgs boson of $B-L$, $q_{\phi}$ is the $B-L$ charge of the field $\phi$, and $\mathcal{L}_{\text{SM}}$ is the standard model Lagrangian. If we ignore the $\theta$ term for $X$, we may assume $y_{N}$ is diagonal and real, $y_{Nij}=\text{diag}(y_{N1},y_{N2},y_{N3})$. For simplicity, we assume all the Yukawa couplings, $y_{Ni}$, are the same order. After $\Phi$ acquires VEV, $\langle\Phi\rangle=v_{B-L}/\sqrt{2}$, the $B-L$ gauge symmetry is broken and the right-handed neutrinos become massive, leaving the unphysical would-be Goldstone boson, $\phi_{M}$, as the $B-L$ charge carrier.
The Yukawa interaction between the right-handed neutrinos and the standard model leptons, Eq. (III.61), equilibrate the $B-L$ charges in the standard model sector and the chiral sector. For simplicity, let us first discuss the case where the $B-L$ charges are totally equilibrated. To estimate the rates, let us take $y_{S}^{2}=y_{S0}^{2}$ with
@@
where $\Delta m_{32}$ is the mass difference between the third and the second generation of the neutrinos, $M_{Ni}$ is the zero-temperature mass of the $i$-th right-handed neutrinos which contribute to the equilibrium process the most, and $v$ is the VEV of the Higgs field. Numerically, $M_{\star}\simeq 6.1\times 10^{14}\,\text{GeV}$. Assuming $y_{S}\lesssim 1$, the most efficient process to equilibrate the $B-L$ charges is scatterings between the top quark and gauge bosons exchanging the Higgs boson and the lepton. Before the symmetry breaking, the rate of the process is $\Gamma_{1}\sim y_{S0}^{2}T$, and for $T\ll\frac{M_{p}}{M_{\star}}M_{Ni}$, the $B-L$ charges are equilibrated, $\Gamma_{1}\gg H(T)$. Therefore, the phase transition temperature of the $B-L$ symmetry breaking, $T_{c}$, must be much smaller than $\frac{M_{p}}{M_{\star}}M_{Ni}$ to meet the first requirement.
@@
On the other hand, the second requirement, decoupling of the interactions before the decay of the magnetic helicity, is almost contradictory to the first requirement. As we have discussed in Sec. II, the decay of the magnetic helicity is expected to be quick after the $B-L$ gauge symmetry is broken. Let us denote the temperature $T_{HD}$. Then, the right-handed neutrino is massive and the scattering rate discussed above is suppressed by the Boltzmann suppression factor, $\exp(-M_{Ni}(T_{HD})/T_{HD})$, where $M_{Ni}(T)=y_{Ni}v_{B-L}(T)/\sqrt{2}$. However, as we have discussed in the previous section, $v_{B-L}(T_{HD})/T_{HD}$ is much smaller than unity, and the scattering rate is not suppressed. Therefore, the $B-L$ charges in the standard model sector and the chiral sector are equilibrated even after the decay of the magnetic helicity. We show the schematic picture of our scenario in this case in Fig. 3.
@@
Figure 3: Schematic picture of our scenario with Majorana neutrinos. Compared with Fig. 2, the washout process does not decouple before the decay of the magnetic helicity and the final $B-L$ charge in the standard model sector is zero.
@@
There can be several possibilities to evade this contradiction. First, we may consider the case where the $B-L$ charges are not totally equilibrated, i.e., the equilibrating interactions are not in the thermal equilibrium. Then, the $B-L$ charges in the standard model sector after the phase transition are a tiny fraction of the initial charges with a negative sign and are not washed out. However, the effects of Eq. (III.60) become stronger as the temperature decreases compared to the Hubble expansion; if the temperature is smaller than around $y_{S0}^{2}M_{p}=M_{Ni}M_{p}/M_{\star}$ well before the right-handed neutrinos decouple, the interaction is in the thermal equilibrium and the final $B-L$ charges in the standard model sector are still washed out. Second, even when the washout process is in the thermal equilibrium, if it decouples within a few Hubble times after the decay of the magnetic helicity, there could be some asymmetry left in the standard model sector. To study this, we need a more detailed analysis of the helicity decay and the washout process near the phase transition. Finally, we may consider the case where the phase transition is of the first order. If the system is, for example, overcooled, the washout process in the Higgs phase can be decoupled due to the Boltzmann suppression factor quickly after the phase transition. This, however, requires a more careful analysis of the helicity decay at the first order phase transition, and we leave this for future work.
@@
If we give up the simplest seesaw model and consider more complicated models, we can generate the baryon asymmetry of the universe along our scenario. In the IR limit, the equilibrating interaction is the dimension-5 Weinberg operator, $\mathcal{O}_{W}=LHLH/\Lambda$, where $\Lambda$ is the cutoff scale. Such a high-dimensional operator is generally stronger in the high-energy scale, but in the simplest seesaw model the operator appears only after the decouple of the right-handed neutrinos, which is the reason why the washout process cannot be decoupled by the time of the helicity decay. Thus, if the dimension-5 operator is generated not by the right-handed neutrinos but by other heavy particles, the washout process can be decoupled earlier and the baryon asymmetry can be generated. For example, we may consider a dimension-6 operator, $\mathcal{O}^{\prime}=LHLH\Phi^{\dagger}/\Lambda^{2}$, which is reduced to the Weinberg operator after the $B-L$ symmetry breaking and assume $y_{S}$ is much smaller than unity and the neutrino masses are mainly generated by this operator. The $B-L$ equilibration is then achieved by $\mathcal{O}^{\prime}$ and the washout process can be decoupled by the helicity decay. Such an operator can be generated by introducing new heavy Dirac fermions, for example.
@@
Up to this point, we have ignored the effect of the flavor symmetry of the standard model particles. In the minimal seesaw model, the right-handed neutrino interactions may break the flavor symmetry for the left-handed leptons but the chirality of the right-handed leptons and the flavor symmetry of the quarks are not broken if the corresponding Yukawa couplings are not in the thermal equilibrium Domcke et al. (2021, 2023). In this case, even though the washout process is strong enough, the chemical potential of the $B-L$ charge and other chiral charges in the standard model sector are not zero before the decay of the magnetic helicity. However, the decay of the magnetic helicity generates the opposite sign of the chemical potentials assuming the washout process is strong enough. Therefore, the final $B-L$ charge in the standard model sector is still zero.

### v2

@@ III Model with Majorana neutrinos
In this section, we examine the dynamics of the BLB-L charge in the fermion sector. We focus on the seesaw model Minkowski (1977); Yanagida (1979); Gell-Mann et al. (1979), in which the interaction between the right-handed neutrinos and the standard model particles results in the standard model neutrinos behaving as Majorana particles at low energies. We investigate if our scenario can account for the generation of the baryon asymmetry of the universe.
@@
Figure 2: Schematic picture of a successful scenario. First, the BLB-L helicity (the blue line) is generated by, for example, the axion inflation, together with the BLB-L charges in the standard model sector (the green line) and the chiral sector (the orange line). Then, the BLB-L charges are (partially) equilibrated by the interactions between these two sectors to wash out the BLB-L charge in the fermion sector. Finally, the BLB-L helicity decays and the BLB-L charges are generated in both sectors. The final BLB-L charge in the standard model sector corresponds to the baryon number of the universe today.
@@
Let us first recapitulate our leptogenesis scenario and summarize the requirements for the model. We show a schematic picture of our scenario in Fig. 2. First, we assume the BLB-L helicity is generated in the early universe. There are several mechanisms to generate the BLB-L helicity, such as the axion inflation Turner and Widrow (1988); Garretson et al. (1992); Anber and Sorbo (2006); Barnaby and Peloso (2011); Barnaby et al. (2011) and the chiral plasma instability Joyce and Shaposhnikov (1997); Akamatsu and Yamamoto (2013); Rogachevskii et al. (2017); Schober et al. (2018); Kamada (2018). As we have discussed in Sec. II, we assume the axion inflation as the generation mechanism in this paper. When the BLB-L helicity is generated, the BLB-L charges in the standard model sector are also generated according to the anomaly equation, Eq. (II.2).******In the case of chiral plasma instability Joyce and Shaposhnikov (1997); Akamatsu and Yamamoto (2013); Rogachevskii et al. (2017); Schober et al. (2018); Kamada (2018), we need to start from a chiral charge for the BLB-L gauge interaction in the fermion sector. The distribution of BLB-L charge in the standard model and chiral sector after BLB-L helicity generation depends on the initial condition. As long as the BLB-L symmetry is gauged, regardless of the detail of the model, some chiral sector is added to cancel the ’t Hooft anomaly of the BLB-L symmetry in the standard model sector. The BLB-L charge is therefore generated in the chiral sector as well. Note that the total BLB-L charge of the universe is zero; the BLB-L charges in the standard model sector and the chiral sector are the same amount with the opposite sign.
@@
As we have discussed in Sec. II, the magnetic helicity of the BLB-L gauge field decays after the BLB-L gauge symmetry is Higgsed. The decay generates the BLB-L charge both in the standard model sector and the chiral sector. If the BLB-L charges in both sectors are conserved separately from the beginning, the final BLB-L charges in both sectors are zero. This can be seen from the integrated form of the anomaly equation, Eq. (II.3); the initial BLB-L charges in both sectors and the magnetic helicity are zero and the final helicity is zero, so that the final BLB-L charges in both sectors are zero.
@@
To generate non-zero BLB-L, we need two requirements. First, we need to introduce the interactions between the standard model particles and the chiral sector to equilibrate, at least partially, the BLB-L charges in both sectors, as explained above. However, if this interaction does not decouple even after the decay of the magnetic helicity, the BLB-L charges generated by the helicity decay in both sectors are also equilibrated, and the final BLB-L charges in both sectors are zero. Thus, secondly, it is also required to decouple the interactions before the decay of the magnetic helicity. Then, the decay of the magnetic helicity generates the difference between the BLB-L charges in the standard model sector and the chiral sector, which remains even today as the baryon number of the universe.
@@
Before discussing the details of the interactions, let us make a few comments on the above summary. Firstly, we have not specified the chiral sector yet. For the seesaw model discussed in this section, the chiral sector is the right-handed neutrinos before the BLB-L symmetry breaking. However, as the right-handed neutrinos become extremely massive after the symmetry breaking, one may wonder what happens after the decoupling of the right-handed neutrinos. Once the right-handed neutrinos become massive, the BLB-L symmetry is spontaneously broken and there is a Goldstone boson, $\phi_{M}$, although the Goldstone boson is eventually eaten by the BLB-L gauge field. The Goldstone boson has the Wess-Zumino-Witten term and cancels the ’t Hooft anomaly Green and Schwarz (1984); Faddeev and Shatashvili (1986); Krasnikov (1985); Babelon et al. (1986); Harada and Tsutsui (1987); Preskill (1991). In other words, the Goldstone boson is chiral and plays the role of the chiral sector. The total BLB-L charge is always zero even today if we include the unphysical contribution from $\phi_{M}$, as BLB-L symmetry is gauged.
@@
Secondly, from the integrated anomaly equation, Eq. (II.3), one may wonder if the BLB-L charge in the standard model sector is always zero given that the initial BLB-L charge and initial and final helicity is vanishing. Contrary to this naive expectation, the final BLB-L charge in the standard model sector is not zero. This is because the interactions equilibrating the BLB-L charges in the standard model sector and the chiral sector explicitly respect only the total BLB-L symmetry and break the BLB-L symmetry in the standard model sector. Schematically, the total derivative of the standard model BLB-L current is modified as
$\partial_{\mu}J^{\mu}_{B-L,\text{SM}}=-\frac{3g_{B-L}^{2}}{16\pi^{2}}X_{\mu\nu}\tilde{X}^{\mu\nu}+\mathcal{O}_{\cancel{B-L}}^{\text{SM}},$ (III.57)
where $\mathcal{O}_{\cancel{B-L}}^{\text{SM}}$ is the operator from the interaction. $\mathcal{O}_{\cancel{B-L}}^{\text{SM}}$ is, in general, not a total derivative and the sum of the BLB-L charge in the standard model sector and the helicity is not conserved anymore. In this sense, $\mathcal{O}_{\cancel{B-L}}^{\text{SM}}$ is the source of the baryon asymmetry of the universe.
@@
$\mathcal{L}_{N\Phi}$ $\equiv\frac{1}{2}y_{Nij}\Phi N_{i}^{T}\mathcal{C}P_{L}N_{j}+\text{H.c.}$ (III.60)
$\mathcal{L}_{Y}$ $\equiv y_{Sij}HN_{i}^{T}\mathcal{C}L_{j}+\text{H.c.}$ (III.61)
@@
where $N_{i}$ is $i$-th ($i=1,2,3$) generation of the right-handed neutrino ††††††Readers should not confuse the name, “right-handed”, with the chirality of the right-handed neutrino. Throughout the paper, we use left-handed Weyl spinors for the chiral fermions., $\Phi$ is the Higgs boson of BLB-L, $q_{\phi}$ is the BLB-L charge of the field $\phi$, and $\mathcal{L}_{\text{SM}}$ is the standard model Lagrangian. If we ignore the $\theta$ term for $X$, we may assume $y_{N}$ is diagonal and real, $y_{Nij}=\text{diag}(y_{N1},y_{N2},y_{N3})$. For simplicity, we assume all the Yukawa couplings, $y_{Ni}$, are the same order. After $\Phi$ acquires VEV, $\langle\Phi\rangle=v_{B-L}/\sqrt{2}$, the BLB-L gauge symmetry is broken and the right-handed neutrinos become massive, leaving the unphysical would-be Goldstone boson, $\phi_{M}$, as the BLB-L charge carrier.
The Yukawa interaction between the right-handed neutrinos and the standard model leptons, Eq. (III.61), equilibrate the BLB-L charges in the standard model sector and the chiral sector. For simplicity, let us first discuss the case where the BLB-L charges are totally equilibrated. To estimate the rates, let us take $y_{S}^{2}=y_{S0}^{2}$ with
@@
where $\Delta m_{32}$ is the mass difference between the third and the second generation of the neutrinos, $M_{Ni}$ is the zero-temperature mass of the $i$-th right-handed neutrinos which contribute to the equilibrium process the most, and $v$ is the VEV of the Higgs field. Numerically, $M_{\star}\simeq 6.1\times 10^{14}\,\text{GeV}$. Assuming $y_{S}\lesssim 1$, the most efficient process to equilibrate the BLB-L charges is scatterings between the top quark and gauge bosons exchanging the Higgs boson and the lepton. Before the symmetry breaking, the rate of the process is $\Gamma_{1}\sim y_{S0}^{2}T$, and for $T\ll\frac{M_{p}}{M_{\star}}M_{Ni}$, the BLB-L charges are equilibrated, $\Gamma_{1}\gg H(T)$. Therefore, the phase transition temperature of the BLB-L symmetry breaking, $T_{c}$, must be much smaller than $\frac{M_{p}}{M_{\star}}M_{Ni}$ to meet the first requirement.
@@
On the other hand, the second requirement, decoupling of the interactions before the decay of the magnetic helicity, is almost contradictory to the first requirement. As we have discussed in Sec. II, the decay of the magnetic helicity is expected to be quick after the BLB-L gauge symmetry is broken. Let us denote the temperature $T_{HD}$. Then, the right-handed neutrino is massive and the scattering rate discussed above is suppressed by the Boltzmann suppression factor, $\exp(-M_{Ni}(T_{HD})/T_{HD})$, where $M_{Ni}(T)=y_{Ni}v_{B-L}(T)/\sqrt{2}$. However, as we have discussed in the previous section, $v_{B-L}(T_{HD})/T_{HD}$ is much smaller than unity, and the scattering rate is not suppressed. Therefore, the BLB-L charges in the standard model sector and the chiral sector are equilibrated even after the decay of the magnetic helicity. We show the schematic picture of our scenario in this case in Fig. 3.
@@
Figure 3: Schematic picture of our scenario with Majorana neutrinos. Compared with Fig. 2, the washout process does not decouple before the decay of the magnetic helicity and the final BLB-L charge in the standard model sector is zero.
@@
There can be several possibilities to evade this contradiction. First, we may consider the case where the BLB-L charges are not totally equilibrated, i.e., the equilibrating interactions are not in the thermal equilibrium. Then, the BLB-L charges in the standard model sector after the phase transition are a tiny fraction of the initial charges with a negative sign and are not washed out. However, the effects of Eq. (III.60) become stronger as the temperature decreases compared to the Hubble expansion; if the temperature is smaller than around $y_{S0}^{2}M_{p}=M_{Ni}M_{p}/M_{\star}$ well before the right-handed neutrinos decouple, the interaction is in the thermal equilibrium and the final BLB-L charges in the standard model sector are still washed out. Second, even when the washout process is in the thermal equilibrium, if it decouples within a few Hubble times after the decay of the magnetic helicity, there could be some asymmetry left in the standard model sector. To study this, we need a more detailed analysis of the helicity decay and the washout process near the phase transition. Finally, we may consider the case where the phase transition is of the first order. If the system is, for example, overcooled, the washout process in the Higgs phase can be decoupled due to the Boltzmann suppression factor quickly after the phase transition. This, however, requires a more careful analysis of the helicity decay at the first order phase transition, and we leave this for future work.
@@
If we give up the simplest seesaw model and consider more complicated models, we can generate the baryon asymmetry of the universe along our scenario. In the IR limit, the equilibrating interaction is the dimension-5 Weinberg operator, $\mathcal{O}_{W}=LHLH/\Lambda$, where $\Lambda$ is the cutoff scale. Such a high-dimensional operator is generally stronger in the high-energy scale, but in the simplest seesaw model the operator appears only after the decouple of the right-handed neutrinos, which is the reason why the washout process cannot be decoupled by the time of the helicity decay. Thus, if the dimension-5 operator is generated not by the right-handed neutrinos but by other heavy particles, the washout process can be decoupled earlier and the baryon asymmetry can be generated. For example, we may consider a dimension-6 operator, $\mathcal{O}^{\prime}=LHLH\Phi^{\dagger}/\Lambda^{2}$, which is reduced to the Weinberg operator after the BLB-L symmetry breaking and assume $y_{S}$ is much smaller than unity and the neutrino masses are mainly generated by this operator. The BLB-L equilibration is then achieved by $\mathcal{O}^{\prime}$ and the washout process can be decoupled by the helicity decay. Such an operator can be generated by introducing new heavy Dirac fermions, for example.
@@
Up to this point, we have ignored the effect of the flavor symmetry of the standard model particles. In the minimal seesaw model, the right-handed neutrino interactions may break the flavor symmetry for the left-handed leptons but the chirality of the right-handed leptons and the flavor symmetry of the quarks are not broken if the corresponding Yukawa couplings are not in the thermal equilibrium Domcke et al. (2021, 2023). In this case, even though the washout process is strong enough, the chemical potential of the BLB-L charge and other chiral charges in the standard model sector are not zero before the decay of the magnetic helicity. However, the decay of the magnetic helicity generates the opposite sign of the chemical potentials assuming the washout process is strong enough. Therefore, the final BLB-L charge in the standard model sector is still zero.